OpenStudy (anonymous):

Gina, Sam and Robby all rented movies from the same video store. They each rented some dramas, comedies and documentaries. Gina rented 11 movies total. Sam rented twice as many dramas, three times as many comedies, and twice as many documentaries as Gina. He rented 27 movies total. If Robby rented 19 movies total with the same number of dramas, twice as many comedies, and twice as many documentaries as Gina, how many movies of each type did Gina rent? 2 dramas, 6 comedies and 3 documentaries 1 dramas, 4 comedies and 6 documentaries 3 dramas, 5 comedies and 3 documentaries 4 dra

OpenStudy (anonymous):

the third answer is correct

OpenStudy (anonymous):

Solve using equations: Gina 11 = a+b+c Sam 27 = 2a+3b+2c Robby 19= a+2b+2c Gina *2 give 22=2a+2b+2c (subtract this from Sam to give) b=5 the subtract Robby from Gina *2 to give a=3

OpenStudy (anonymous):

i used guess and check XD but econmodeller is correct there are many ways to do this type of problem as well